a) (i) Use EDTA structure to describe hexadentate chelating agent (2mks) (ii) State four types of EDTA titrations (2mks) b)ii)Consider the titration of 50mL Mg 2+ sample in a pH = 10.00 buffer with 0.05M EDTA as titrant. Calculate the amount of Mg in ppm if the titre is 5ml of EDTA (3mks) ii) Calculate the pH of a 1000-mL buffer solution containing 0.200M NH4OH and 0.300M NH4CI. (3mks)
1
Expert's answer
2020-04-08T13:37:15-0400

a) I. A hexadentate ligand in coordination chemistry is a ligand that combines with a central metal atom with six bonds. (Picture: generic structure of a metal complex of hexadentate EDTA)


II. Acid-base titrations

Redox titrations

Complexometric titrations

Precipitation titrations

b) I. CMg = 24 × 0.05 / 10 × 103 = 120 ppm

II. pH = pKa + lg(aNH4OH / aNH4Cl) = 9.25 + lg(0.2 / 0.3) = 8.67